OpenStudy (anonymous):

y=3(w-y) solve for y

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

y=3(w-y) <-- distribute the 3 y=3w-3y +3y +3y ___________ 4y=3w y=3w/4

OpenStudy (anonymous):

thank u so much and grat explication lol :)

OpenStudy (anonymous):

haha no problem